# Computing Balanced Solutions for Large International Kidney Exchange Schemes When Cycle Length Is Unbounded

### v1

@@ 4 Simulations
Initial allocations. For the initial allocations $y$ we use the Banzhaf value, Shapley value, nucleolus, benefit value and contribution value. The Shapley value $\phi(N,v)$ [36] is defined by $\phi_{p}(N,v)=\sum_{S\subseteq N\backslash\{p\}}\frac{|S|!(n-|S|-1)!}{n!}(v(S\cup\{p\})-v(S))$. To define the next solution concept, we first introduce unnormalized Banzhaf value $\psi_{p}(N,v)$ [30] defined by $\psi_{p}(N,v):=\sum_{S\subseteq N\backslash\{p\}}\frac{1}{2^{n-1}}(v(S\cup\{p\})-v(S))$. As $\psi_{p}$ may not be an allocation, the (normalized) Banzhaf value $\overline{\psi}_{p}(N,v)$ of a game $(N,v)$ was introduced and defined by $\overline{\psi}_{p}(N,v):=\frac{\psi_{p}(N,v)}{\sum_{q\in N}\psi_{q}(N,v)}\cdot v(N)$. Whenever we mention the Banzhaf value, we will mean $\overline{\psi}(N,v)$. For the Shapley value and Banzhaf value, we were still able to implement a brute force approach relying on the above definitions.
@@
We now define the nucleolus. The excess for an allocation $x$ of a game $(N,v)$ and a non-empty coalition $S\subsetneq N$ is defined as $e(S,x):=x(S)-v(S)$. Ordering the $22^{n}-$ 2 excesses in a non-decreasing sequence yields excess vector $e(x)\in\mathbb{R}^{2^{n}-2}$. The nucleolus of a game ($N,v$) is the unique allocation [35] that lexicographically maximizes $e(x)$ over the set of allocations $x$ with $x_{i}\geq v(\{i\})$ (assuming this set is nonempty, as holds in our case). To compute the nucleolus, we use the Lexicographical Descent method of [10], which is the state-of-the-art method in nucleolus computation.
@@
The surplus of a game $(N,v)$ is $\mbox{surp}=v(N)-\sum_{p\in N}v(\{p\})$. One can allocate $v(\{p\})+\alpha_{p}\cdot\mbox{surp}$, as long as $\sum_{p\in N}\alpha_{p}=1$, to a player $p\in N$. If $\alpha_{p}=\frac{v(N)-v(N\setminus\{p\})-v(\{p\})}{\sum_{p\in N}(v(N)-v(N\setminus\{p\})-v(\{p\}))}$, we obtain the benefit value. If $\alpha_{p}=\frac{v(N)-v(N\setminus\{p\})}{\sum_{p\in N}(v(N)-v(N\setminus\{p\}))}$, we get the contribution value. Both values are easy to compute, but may not exist if the denominator is zero, which did not happen in our simulations.
Optimal solutions. We compute weakly and strongly close optimal solutions by solving a sequence of ILPs, as described in Section 3. We used g++ version 11 on Ubuntu 20.04 for our C++ implementation. The values $v(S)$ of the partitioned permutation games used in the ILPs are obtained by solving a maximum weight perfect matching problem (see Section 1), for which we use the package of [24].
@@
Computational environment and scale. We run our simulations both without and with (“$+c$”) using the credit system, and for the settings, where an arbitrary optimal solution (“arbitrary”), weakly optimal solution (“$d_{1}$”) or strongly optimal optimal solution (“lexmin”) is chosen. This leads to the following five scenarios for each of the five selected solution concepts, the Shapley value, nucleolus, Banzhaf value, benefit value and contribution value: arbitary, d1, d1+c, lexmin, and lexmin+c. Note that for the arbitrary scenario, the use of credits is irrelevant. Hence, in total, we run the same set of simulations for $5\times 5=25$ different combinations of scenarios and initial allocations. As we have seven different country sizes $n$ and 100 initial compatibility graphs $G_{i}$, our total number of 24-round simulation instances is $25\times 7\times 100=17500$.
All simulations were run on a dual socket server with Xeon Gold 6238R processors with 2.2 GHz base speed and 512GB of RAM, where each simulation was given eight cores and 16GB of RAM.
@@
Evaluation measures. Let $y^{*}$ be the total target allocation of a single simulation instance, i.e., $y^{*}$ is obtained by taking the sum of the 24 initial allocations of each of the 24 rounds. Let ${\mathcal{C}}^{*}$ be the union of the chosen maximum cycle packings in each of the 24 rounds. We use the total relative deviation defined as $\frac{\sum_{p\in N}|y_{p}^{*}-s_{p}({\mathcal{C}^{*}})|}{|{\mathcal{C}^{*}}|}$. For each choice of initial allocation and choice of scenario, we run 100 instances. We take the average of the 100 relative total deviations to obtain the average total relative deviation. Taking the maximum relative deviation $\frac{\max_{p\in N}|y_{p}^{*}-s_{p}({\mathcal{C}^{*}})|}{{\mathcal{C}^{*}}}$ gives us the average maximum relative deviation as our second evaluation measure. As we shall see, this evaluation measure leads to the same conclusions.
@@
Simulation Results for ${\mathbf{\ell=\infty}}$ and Comparison with ${\mathbf{\ell=2}}$. In Figure 4 we display our main results. In this figure, we compare different solutions concepts under different scenarios for $\ell=\infty$. The figure also shows the effects of weakly and strongly close solutions and the credit system. As solution concepts have different complexities, we believe such a comparison might be helpful for policy makers for choosing a solution concept and scenario. As expected, using an arbitrary maximum cycle packing in each round makes the kidney exchange scheme significantly more unbalanced, with average total relative deviations over 4% for all initial allocations $y$. The effect of both selecting a strongly close solution (to ensure being close to a target allocation) and using a credit function (for fairness, to keep deviations small) is significant.
The above observations are in line with the results under the setting where $\ell=2$ [8]. However, for $\ell=2$, the effect of using arbitrary optimal solutions is worse, while deviations are smaller than for $\ell=\infty$ when weakly close or strongly close optimal solutions are chosen. From Figure 4 we see that the Shapley value and the Banzhaf value in the lexmin+c scenario provide the smallest deviations from the target allocations (just like when $\ell=2$ [8]). However, differences are small and, as mentioned, which solution concept to select is up to the policy makers of the IKEP. Moreover, from Figure 4, we can also see the benefits of $\ell=\infty$ over $\ell=2$ for the Shapley value.
Figure 5 shows that if we use the average maximum relative deviation instead of the average total relative deviation, then we can see the same pattern as in Figure 4.
From Figure 6 we see that when $\ell=\infty$ instead of $\ell=2$, it is possible to achieve 46% more kidney transplants. The $\ell=\infty$ setting is not realistic. Namely, Figure 7 show that long cycles, with even more than 400 vertices, may occur. Figure 7 also shows that these long cycles all happen in the first round. We note that before our experiments, the relation between increase in transplants versus increase in cycle length had not been researched by simulations. Finding out about this was our main motivation for our simulations.
@@
Figure 4: Average total relative deviations for each of the five solution concepts under the five different scenarios, where the number of countries $n$ is ranging from $4$ to $10$. For comparison, the bottom figure displays a result from [8] for $\ell=2$, namely for the Shapley value, which behaved best for $\ell=2$.
@@
Figure 5: Average maximum relative deviations for each of the five solution concepts under the five different scenarios, where the number of countries $n$ is ranging from $4$ to $10$.
Figure 6: The total number of kidney transplants when the Shapley value is chosen, both when $\ell=2$ (five scenarios, with overlapping plots) and when $\ell=\infty$ (five scenarios, with overlapping plots). The number of countries $n$ is ranging from $4$ to $10$.
@@
Figure 7: Cycle distribution for the Shapley value in the lexmin+c scenario, where the x-axis represents the cycle length, and the y-axis shows the number of kidney transplants involved in a cycle of that length.

### v2

@@ 4.2 The Initial Allocations
Whenever we mention the Banzhaf value, we will mean $\overline{\psi}(N,v)$. To compute the Shapley value and Banzhaf value in our simulations for even up to $n=10$ countries, we were still able to implement a brute force approach relying on the above definitions.
@@
We now define the nucleolus. The excess for an allocation $x$ of a game $(N,v)$ and a nonempty coalition $S\subsetneq N$ is defined as $e(S,x):=x(S)-v(S)$. Ordering the $22^{n}-$ 2 excesses in a non-decreasing sequence yields excess vector $e(x)\in\mathbb{R}^{2^{n}-2}$. The nucleolus of a game ($N,v$) is the unique allocation [36] that lexicographically maximizes $e(x)$ over the set of allocations $x$ with $x_{i}\geq v(\{i\})$ (assuming this set is nonempty, as is the case for partitioned permutation games). To compute the nucleolus, we use the Lexicographical Descent method of [11], which is the state-of-the-art method in nucleolus computation.
Next, we define the tau value, introduced by Tijs [41]. First, let $b_{p}=v(N)-v(N\setminus p)$ be the utopia payoff of $p\in N$, leading to a vector of utopia payoffs $b\in\mathbb{R}^{N}$. Now, for $p\in S\subseteq N$, we let
$R(S,p)=v(S)-\sum_{q\in S\setminus p}b_{q}$
be the remainder for player $p$ in $S$, which is what would remain if all players apart from $p$ would leave a coalition $S$ with their utopia payoff. We now set for $p\in N$,
$a_{p}=\max_{S\ni p}R(S,p),$
leading to the vector $a\in\mathbb{R}^{N}$ of minimal rights. A game $(N,v)$ is quasibalanced if $a\leq b$ and $a(N)\leq v(N)\leq b(N)$. For a quasibalanced game, the tau value $\tau$ of a game $(N,v)$ is defined by setting for $p\in N$,
$\tau_{p}=\gamma a_{p}+(1-\gamma)b_{p},$
where $\gamma\in[0,1]$ is determined by $\tau(N)=v(N)$. Note that $\gamma$ is unique, unless $a=b$ (in which case we can simply take $\tau=b$).
The tau value is only defined for quasibalanced games. However, all partitioned permutation games have a nonempty core by Theorem 1.3, and thus they are quasibalanced. As for the Shapley value and Banzhaf value, we were still able to compute the tau value by using a brute force approach based on its definition even for $n=10$ countries.
@@
The surplus of a game $(N,v)$ is defined as $\mbox{surp}=v(N)-\sum_{p\in N}v(\{p\})$. As long as $\sum_{p\in N}\alpha_{p}=1$, we can allocate $v(\{p\})+\alpha_{p}\cdot\mbox{surp}$ to each player $p\in N$. In this way we obtain the benefit value by setting for each $p\in N$,
$\alpha_{p}=\frac{v(N)-v(N\setminus\{p\})-v(\{p\})}{\sum_{p\in N}(v(N)-v(N\setminus\{p\})-v(\{p\}))},$
and we obtain the contribution value by setting for each $p\in N$,
$\alpha_{p}=\frac{v(N)-v(N\setminus\{p\})}{\sum_{p\in N}(v(N)-v(N\setminus\{p\}))}.$
Both the benefit value and contribution value are easy to compute, as we only need to compute $12n-$ 1 $v$-values. We also note that that the benefit value and contribution value do not exist if the denominator is zero. However, this did not happen in our simulations.
Finally, we consider a recent variant of the Banzhaf value, which was introduced in [10]. We first define for some round $h$ in an IKEP the credit-adjusted game $(N^{h},\overline{v}^{h})$, in which the credits are directly incorporated, that is, $\overline{v}^{h}(S)=v^{h}(S)+\sum_{p\in S}c^{h}_{p}$ for every $S\subseteq N$. As explained in [10], this alternative way of processing credits does not make any difference for any of the solution concepts that we consider except for one: the (normalized) Banzhaf value. We denote the adjusted Banzhaf value as the Banzhaf* value. In our simulations, we compute it by using brute force as we do for the Banzhaf value.
4.3 The Optimal Solutions
As optimal solutions we compute the following for each computability graph:
  • 1.
    an arbitrary optimal solution;
  • 2.
    a weakly close optimal solution; and
  • 3.
    a strongly close optimal solution.
In this way we can measure the effect of using weakly close optimal solutions over arbitrarily chosen ones as well as the effect of using strongly close optimal solutions over weakly close ones. We compute the weakly and strongly close optimal solutions by solving a sequence of ILPs, as described in Section 3. We use Gurobi Optimizer version 10.0.02(linux64) to create environments, build and solve ILPs (via the Gurobi C++ interface on a cloud system). As not every ILP might be solved within a reasonable time, we give each ILP a time limit of one hour. We recall that the values $v(S)$ of the partitioned permutation games used in the ILPs have already been computed for the initial allocations, and we simply reuse them.
4.4 Computational Environment and Scale
@@
In our large-scale experimental study, we run our simulations both without and with (“$+c$”) the credit system, and as mentioned for the settings, where an arbitrary optimal solution (“arbitrary”), weakly optimal solution (“$d_{1}$”) or strongly close optimal solution (“lexmin”) is chosen. This leads to the following five scenarios for each of the six selected solution concepts, the Shapley value, Banzhaf value, nucleolus, tau value, benefit value and contribution value:
  • arbitary,
  • d1,
  • d1+c,
  • lexmin,
  • lexmin+c;
plus two scenarios for the Banzhaf* value:
  • d1+c,
  • lexmin+c,
as without credits the lhe Banzhaf* value coincides with the Banzhaf value.
Note that for the arbitrary scenario, the use of credits is irrelevant. Hence, in total, we run the same set of simulations for $5\times 6+2=32$ different combinations of scenarios and solution concepts. We consider both the situation where all countries have the same size and a situation where they have different sizes. Moreover, we have seven different country sizes $n$ and 100 compatibility graphs $G_{i}$. Hence, the scale of our experimental study is very large. Namely, the total number of 24-round simulation instances is
$32\times 2\times 7\times 100=44800$
All simulations were run on a dual socket server with AMD EPYC 7702 64-Core Processor with 2.00 GHz base speed and 256GB of RAM, where each simulation was given eight cores and 1G temporary disk space. As we will compare our results for $\ell=\infty$ with the known results for $\ell=2$ from [10], let us point out that the simulations in [10] were run on a desktop PC with AMD Ryzen 9 5950X 3.4 GHz CPU and 128 GB of RAM, running on Windows 10 OS and C++ implementation in Visual Studio (the code is available in GitHub repository [6]).
4.5 Evaluation Measures
@@
Let $y^{*}$ be the total initial allocation of a single simulation instance, that is, $y^{*}$ is obtained by taking the sum of the 24 initial allocations of each of the 24 rounds of that instance. Let ${\mathcal{C}}^{*}$ be the union of the chosen maximum cycle packings in each of the 24 rounds. We use the total relative deviation defined as
$\frac{\sum_{p\in N}|y_{p}^{*}-s_{p}({\mathcal{C}^{*}})|}{|{\mathcal{C}^{*}}|}.$
For each choice of solution concept, choice of scenario and choice of number of countries, we run 100 simulation instances. We take the average of the 100 total relative deviations to obtain the average total relative deviation. Taking the maximum relative deviation
$\frac{\max_{p\in N}|y_{p}^{*}-s_{p}({\mathcal{C}^{*}})|}{|{\mathcal{C}^{*}}|}$
gives us the average maximum relative deviation as our second evaluation measure. As we shall see, both evaluation measures lead to the same conclusions.
4.6 Simulation Results for ${\mathbf{\ell=\infty}}$ and Comparison with ${\mathbf{\ell=2}}$
@@
In Figure 4 we display our main results for equal country sizes, that is, when all countries are of the same size (see also Section 4.1). In Figure 4, we compare different solution concepts under different scenarios for $\ell=\infty$. Figure 4 also shows the effects of weakly and strongly close solutions and the credit system. As solution concepts have different complexities, we believe such a comparison might be helpful for policy makers in choosing a specific solution concept and scenario. As expected, using an arbitrary maximum cycle packing in each round makes the kidney exchange scheme significantly more unbalanced, with average total relative deviations over 4% for all initial allocations $y$. The effect of both selecting a strongly close solution (to ensure being close to a target allocation) and using a credit function (for fairness, to keep deviations small) is significant.
@@
Figure 4: Average total relative deviations for each of the seven solution concepts under the five different scenarios for equal country sizes, where the number of countries $n$ is ranging from $4$ to $10$. For comparison, the lower right figure displays a result from [10] for $\ell=2$, namely for the Banzhaf* value, which behaved best for $\ell=2$. We recall that the Banzhaf value and Banzhaf* value coincide when credits are not incorporated, and this is also reflected in the two corresponding figures.
The above observations are in line with the results under the setting where $\ell=2$ [10]. However, for $\ell=2$, the effect of using arbitrary optimal solutions is much worse, while deviations are smaller than for $\ell=\infty$ when weakly close or strongly close optimal solutions are chosen. From Figure 4 we see that the Banzhaf* value in the lexmin+c scenario provide the smallest deviations from the target allocations (as when $\ell=2$ [10]). However, all solution concepts are within 1.52% (for lexmin+c) and, as mentioned, which solution concept to select should be decided by the policy makers of the IKEP.
@@
Figure 5: Average total relative deviations for each of the seven solution concepts under the five different scenarios for varying country sizes, where the number of countries $n$ is ranging from $4$ to $10$. For comparison, the lower right figure displays a result from [10] for $\ell=2$, namely for the Banzhaf* value, which behaved best for $\ell=2$.
Figure 5 shows the same kind of results as Figure 4 but now for varying country sizes, that is, when we have small, medium and large countries, divided exactly as in [10] for $\ell=2$ (see also Section 4.1). Note that subject to minor fluctuations we can draw the same conclusions from Figure 5 for varying country sizes as we did from Figure 4 for equal country sizes.
Figures 4 and 5 highlight the comparison between different scenarios. For an easier comparison between the effects of choosing different solution concepts for prescribing the initial allocations, we grouped together all the lexmin+c plots in of Figures 4 and 5 in Figure 6 and all the d1+c plots from Figures 4 and 5 in Figure 7. Figures 6 and 7 both show an ordering from the Banzhaf* value, which has the best performance, to the contribution value which has the worst in almost all cases for the two most important scenarios lexmin+c and d1+c.
We note that Figures 6 and 7 also show that (as expected) the effect of varying the country sizes is stronger if the number $n$ of countries is relatively small, especially when $n\in\{4,5,6\}$.
Our Second Evaluation Measure. We can draw exactly the same conclusions as above if we use the average maximum relative deviation instead of the average total relative deviation. We refer to Appendix 0.A for the analogs of Figures 57 if we use the average maximum relative deviation as our evaluation measure.
Figure 6: Average total relative deviations for all solution concepts in the lexmin+c scenario, where the number of countries $n$ ranges from $4$ to $10$.
Figure 7: Average total relative deviations for all solution concepts in the d1+c scenario, where the number of countries $n$ ranges from $4$ to $10$.
Incomplete Instances. In the simulations for $\ell=2$ [10], ILPs were not used, and all simulation instances were solved to optimality. However, our new simulations for $\ell=\infty$ heavily relied on ILPs, and we recall that we imposed a time limit of one hour for our ILP solver (Gurobi) to solve an ILP. Given the nature of ILPs, it is not surprising that there were several ILPs that could not be solved within the $1$-hour time limit. We call the corresponding simulation instances incomplete. So, for such instances, there was one ILP in some round, which our ILP solver could not handle within the $1$-hour time limit. We call call other simulation instances complete.
All ILPs that were not finished within one hour were of the ($\text{ILP}_{d_{t}}$) type, almost always for $t>1$. In fact, for equal country sizes, only one single incomplete simulation instance occurred at d1 / d1+c (for nine countries, using the tau value). Tables 1 and 2 summarize the distribution of incomplete simulation instances aggregated over different choices of solution concepts or scenarios. We refer to Appendix 0.B for a complete breakdown of the averages in both these tables.
Given that the solutions found for incomplete simulation instances are all maximum cycle packings (which might not be weakly or strongly close), policy makers could therefore still decide to use them. For this reason, we decided to include the incomplete simulation instances in Figures 48. We believe this was justified after doing some additional research. That is, we also constructed the same figures as Figures 4 and 5 but without the incomplete simulation instances; see Appendix 0.B. It turned out that the largest percentage points difference in Figure 4 of the average total deviations is only 0.059% (with an average of 0.0024%). Hence, the quality of the “current-best” optimal solutions for the incomplete simulation instances are almost indistinguishably close to those for the complete simulation instances.
#unsolved / n 4 5 6 7 8 9 10 Total
Shapley value 0.00% 0.00% 0.00% 1.50% 1.75% 6.75% 3.50% 1.93%
Nucleolus 0.00% 0.00% 0.00% 0.50% 0.00% 0.00% 0.25% 0.11%
Benefit value 0.25% 0.25% 1.25% 0.25% 0.75% 0.75% 0.00% 0.50%
Contribution value 0.75% 0.75% 1.75% 1.00% 1.50% 0.50% 2.25% 1.21%
Banzhaf value 0.00% 0.75% 1.50% 1.75% 0.50% 4.25% 4.50% 1.89%
Tau value 0.00% 0.25% 0.00% 0.00% 0.00% 1.00% 0.00% 0.18%
Banzhaf* value 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.50% 0.07%
total: d1 0.00% 0.00% 0.00% 0.00% 0.00% 0.17% 0.00% 0.02%
total: lexmin 0.17% 0.33% 0.50% 1.33% 0.83% 4.17% 3.33% 1.52%
total: d1+c 0.00% 0.00% 0.00% 0.00% 0.00% 0.14% 0.00% 0.02%
total: lexmin+c 0.43% 0.86% 2.14% 1.71% 1.86% 3.71% 3.29% 2.00%
Table 1: Average number of incomplete simulation instances for equal country sizes.
#unfinished / n 4 5 6 7 8 9 10 Total
Shapley value 0.00% 0.00% 0.00% 0.50% 0.00% 0.50% 2.00% 0.43%
Nucleolus 0.00% 0.00% 0.00% 0.00% 0.00% 0.75% 0.00% 0.11%
Benefit value 0.25% 0.00% 0.75% 0.00% 0.00% 0.25% 0.75% 0.29%
Contribution value 0.00% 0.00% 0.50% 0.00% 0.50% 0.50% 0.25% 0.25%
Banzhaf value 0.00% 1.00% 0.25% 0.25% 1.25% 0.50% 2.00% 0.75%
Tau value 0.00% 0.00% 0.00% 0.00% 0.25% 0.00% 0.00% 0.04%
Banzhaf* value 0.00% 1.00% 0.00% 0.00% 0.50% 0.50% 2.00% 0.57%
total: d1 0.00% 0.17% 0.00% 0.00% 0.00% 0.17% 0.00% 0.05%
total: lexmin 0.00% 0.17% 0.33% 0.33% 0.50% 0.50% 1.67% 0.50%
total: d1+c 0.00% 0.29% 0.14% 0.00% 0.14% 0.00% 0.00% 0.08%
total: lexmin+c 0.14% 0.29% 0.43% 0.14% 0.71% 1.00% 2.00% 0.67%
Table 2: Average number of incomplete simulation instances for varying country sizes.
Number of Kidney Transplants versus Cycle Length. When $\ell=\infty$ instead of $\ell=2$, we expect more kidney transplants as exchange cycles may now have any size. In Figures 8 and 9 we quantified this. From Figure 8 we see that when $\ell=\infty$ instead of $\ell=2$, it is possible to achieve 46% more kidney transplants. This is irrespective of the chosen scenario, as the total number of kidney transplants are nearly identical throughout the five scenarios (as for $\ell=2$). The $\ell=\infty$ setting is not realistic of course. Namely, Figure 9 show that long cycles, with even more than 400 vertices, may occur. Figure 9 also shows that these long cycles all happen in the first round. We note that before our experiments, the relation between increase in transplants versus increase in cycle length had not been researched by any simulations. Finding out about this relation through an extensive set of simulations was one of our main motivations.
Figure 8: Left: the total number of kidney transplants when the Banzhaf* value is chosen when the countries are of the same size. The dotted lines correspond to the case where $\ell=\infty$ and the solid lines are taken from [10] and correspond to the case where $\ell=2$. Right: the same figure but now for the case where the countries have different sizes. Note that in both figures some of the plots are overlapping.
@@
Figure 9: Cycle distribution for the Banzhaf* value in the lexmin+c scenario, where the x-axis represents the cycle length, and the y-axis shows the number of kidney transplants involved in a cycle of that length.
Influence of Using Credits. We now discuss the power of our credit system. Does it effectively reduce deviations from the target allocation? Can it keep deviations on a consistently low level in the long-term? Theoretically it is possible that credits keep accumulating, which would make them ineffective (see [10] for a theoretical example of this behaviour when $\ell=2$).
In Figures 10 and 11 we show to what extent credits accumulate for equal and varying country sizes, respectively. If credits are not incorporated, we can still compute and track them as we did in these two figures. We note that over a period of 24 rounds, credits accumulate more and more (but at different rates) under all three scenarios arbitary, d1 and lexmin, even though especially under d1 and lexmin we do find solutions that are relatively close to the target allocations (as we saw from Figures 4 and 5). However, only when we also incorporate credits we not only find solutions that are close to the target solutions but that also ensure stability. We find that the Banzhaf* value under lexmin+c is also the best in maintaining consistently low levels of credits, much like in the case of $\ell=2$ [10]. Again, the main difference between $\ell=\infty$ and $\ell=2$ is that the arbitrary scenario performs much worse for $\ell=2$.
Figure 10: Average credits for each of the seven solution concepts under the five different scenarios for equal country sizes, where the number of countries is $n=10$ and the period ranges from $1$ to $24$. For comparison, the lower right figure displays a result from [10] for $\ell=2$, namely for the Banzhaf* value, which also behaved best for $\ell=2$.
Figure 11: Average credits for each of the seven solution concepts under the five different scenarios for varying country sizes, where the number of countries is $n=10$ and the period ranges from $1$ to $24$. For comparison, the lower right figure displays a result from [10] for $\ell=2$, namely for the Banzhaf* value, which also behaved best for $\ell=2$.
Deviations for Worst-off Countries. Finally, we consider the ratio of the average maximum relative deviation and the average total relative deviation as a measure for the proportion of the total country deviations being due to the worst-off country. We call this ratio the relative ratio. For stability reasons, low relative ratios are preferable. From Table 3 we conclude that for the most interesting scenario lexmin+c and equal country sizes, the relative ratio decreases as the number of countries increases (as expected). In this figure we also included the relative ratios for $\ell=2$ (generated from the data used in [10]). We note that the relative ratio is slightly better for $\ell=2$. This extends to varying countries as well; see Appendix 0.C. In this appendix we also included the figures for the other scenarios as well. From these figures, we note that for d1+c the relative ratio for $\ell=\infty$ is lower than for $\ell=2$. The arbitrary scenario showed consistently lower deviations for $\ell=\infty$ than for $\ell=2$. However, this is not the case for the relative ratio. For equal country sizes, $\ell=\infty$ leads to consistently higher relative ratios than $\ell=2$ does, whereas for varying country set sizes the opposite holds. For the remaining scenarios (lexmin and d1), there is no clear distinction.
n 4 5 6 7 8 9 10
Shapley value $\ell=\infty$ 46.82% 39.70% 38.99% 34.55% 31.40% 31.41% 28.17%
$\ell=2$ 41.70% 37.92% 35.29% 30.85% 28.00% 26.71% 26.52%
difference 5.12% 1.78% 3.70% 3.70% 3.40% 4.69% 1.66%
Banzhaf value $\ell=\infty$ 45.91% 40.92% 39.50% 36.87% 30.95% 30.57% 27.61%
$\ell=2$ 43.90% 37.07% 34.88% 30.05% 29.27% 26.89% 26.26%
difference 2.01% 3.85% 4.62% 6.82% 1.68% 3.68% 1.35%
nucleolus $\ell=\infty$ 47.35% 39.50% 38.25% 35.21% 32.95% 29.81% 27.02%
$\ell=2$ 45.85% 40.05% 36.57% 33.68% 30.84% 29.44% 26.75%
difference 1.51% -0.54% 1.67% 1.52% 2.11% 0.36% 0.27%
tau value $\ell=\infty$ 47.11% 41.16% 39.47% 34.02% 33.91% 31.57% 27.07%
$\ell=2$ 42.00% 38.53% 37.70% 32.41% 30.31% 28.68% 27.69%
difference 5.11% 2.63% 1.78% 1.62% 3.60% 2.89% -0.62%
benefit value $\ell=\infty$ 46.24% 40.37% 39.72% 34.89% 34.08% 30.44% 27.10%
$\ell=2$ 42.91% 38.13% 36.62% 30.11% 31.76% 28.57% 26.02%
difference 3.33% 2.25% 3.11% 4.78% 2.32% 1.87% 1.08%
contribution value $\ell=\infty$ 44.41% 37.06% 38.30% 34.77% 31.76% 28.55% 26.41%
$\ell=2$ 43.19% 40.15% 37.50% 31.94% 29.80% 28.19% 25.90%
difference 1.23% -3.09% 0.80% 2.83% 1.96% 0.36% 0.51%
Banzhaf* value $\ell=\infty$ 45.12% 40.51% 38.89% 35.16% 31.22% 29.59% 27.22%
$\ell=2$ 43.33% 37.57% 34.47% 30.42% 28.48% 27.23% 25.33%
difference 1.78% 2.94% 4.42% 4.74% 2.74% 2.36% 1.89%
Table 3: Relative ratios for equal country sizes under lexmin+c, for $\ell=\infty$, $\ell=2$ and their difference.
4.7 Computation Times
Table 4 summarizes the average computational time of solving a single (24-round) simulation instance. As expected, computing initial allocations using the two easy-to-compute solution concepts (benefit and contribution values) is inexpensive, especially compared to the other, more sophisticated but hard-to-compute solution concepts.
In Table 4 we also see that the computation time for the arbitrary scenario, which does not require initial allocations, is less time-consuming than for the other four scenarios. This is in line with Theorem 1.1, which gave us a polynomial time algorithm for computing an arbitrary optimal solution, whereas we needed to rely on using ILPs for computing optimal solutions that are weakly close (d1 and d1+c) or even strongly close (lexmin and lexmin+c). From Table 4, it can be clearly seen that the computation of the ILP series requires the majority of the work, especially for the lexmin and lexmin+c scenarios: these involve longer series of ILPs than d1 and d1+c, for which we need to solve only ($\text{ILP}_{d_{1}}$).
On a side note, we observe from Table 4 that when $n$ increases, the computation time increases roughly within the expected rate, apart from a few notable exceptions: Gurobi’s ILP solver handles, for some reason, ($\text{ILP}_{d_{1}}$) far more easily for $n=6$ than for $n=5$ (or even $n=4$). We have no explanation for this; it may well be due to the nature of Gurobi, in which we have no insights.
CPU time / n 4 5 6 7 8 9 10
data preparation 0.01 0.01 0.01 0.01 0.01 0.01 0.01
graph building 0.04 0.04 0.04 0.04 0.04 0.04 0.04
Shapley value 12.32 25.28 48.32 94.70 188.7 393 745.2
Nucleolus 10.75 22.33 44.41 86.83 165.5 342 707.8
Benefit value 7.87 8.41 9.64 11.31 12.68 14.45 15.81
Contribution value 7.39 8.80 10.35 12.44 14.15 16.68 17.00
Banzhaf value 10.56 21.73 42.06 89.00 174.75 343.06 689.62
Banzhaf* value 13.03 25.48 50.98 89.12 173.76 359.56 687.96
Tau value 17.76 29.94 56.70 99.66 199 358.1 796.2
total: arbitrary 11.63 20.35 35.35 65.49 123.81 257.33 530.78
total: d1 66.81 85.85 42.97 127.49 171.84 315.08 473.38
total: lexmin 305.04 688.20 1044.89 1929.32 2771.54 4912.33 6737.04
total: d1+c 93.26 87.09 43.93 121.11 166.00 330.56 504.52
total: lexmin+c 470.29 1030.02 2010.87 2514.62 4032.30 5764.23 7990.36
Table 4: Average CPU time for a single 24-round simulation instance, broken down into the different computational tasks. Here, the times for data preparation and graph building are average times taken over all scenarios and solution concepts, whereas the time for each solution concept is the average time taken over all scenarios. The total times for the scenarios are average times taken over all solutions concepts, where “total” refers to the total computation time, which includes computing the initial allocations, and only excludes the time for data preparation and graph building.